Jack Hemmings bought a 3-month British pound futures contract for $1.4400 British pound only to see the dollar appreciate to a value of $1.4250 at which he sold the pound futures. If each pound futures is for an amount of British pound 62,500, how much money did jack gain or lose from his speculation with pound futures?

Alcoa has a DKr 3 million receivable due in 6 months. What value can Alcoa lock-in for its receivable if it executes a forward market hedge today when the forward rate is $.01729?

Alcoa has DKr 3 million receivable due in 6 months. The following information on Dkr options are available: 6-month put option with a strike price of $.1745 is priced at $0.004 and a 6-month call option with a strike price of $0.1750 is priced at $0.005. If Alcoa executes options to hedge, what value for receivables it can lock in?
Answer #1
JACK HEMMINGS
Loss = (Purchase price-sale price)*pounds per contract
= (1.44-1.425)*62500
= 937.5 Pounds
Alcoa
a) Value locked by entering forward hedge = 3,000,000*0.01729
= 51870 dollars
b) Since alcoa has DKr receivable she is afraid of DKr falling & USD appreciating
Hence to hedge ,she will purchase call option on dollars
Price locked = (Strike price* amount receivble)-premium paid
= (0.1745*3,000,000)-{0.004*(3,000,000*0.1745)}
= $523500-$2094
= 521406 dollars